# Data Taggants: Dataset Ownership Verification via Harmless Targeted Data Poisoning

### v1

@@ Abstract
Dataset ownership verification, the process of determining if a dataset is used in a model’s training data, is necessary for detecting unauthorized data usage and data contamination. Existing approaches, such as backdoor watermarking, rely on inducing a detectable behavior into the trained model and on a part of the data distribution. However, these approaches have limitations, as they can be harmful to the model’s performance, require unpractical access to the model’s internals. Furthermore, previous approaches lack guarantee against false positives.
This paper introduces data taggants, a novel non-backdoor dataset ownership verification technique, specifically designed for image classification datasets. Our method uses pairs of out-of-distribution samples and random labels as secret keys, and leverages clean-label targeted data poisoning to subtly alter a dataset so that models trained on this dataset respond to the key samples with the corresponding key labels. The keys are built as to allow for statistical certificates with black-box access only to the model.
We validate our approach through comprehensive experiments on ImageNet1k using ViT and ResNet models with state-of-the-art training recipes. Our findings demonstrate that data taggants can reliably identify models trained on the protected dataset with high confidence, without compromising validation accuracy. We show the superiority of our approach over backdoor watermarking in terms of performance and stealthiness. Furthermore, our method exhibits robustness against various defense mechanisms from data poisoning literature, accross different architectures, and dataset modifications.
@@ 1 Introduction
Dataset ownership verification approaches aim to equip dataset owners with the ability to track the usage of their data in specific trained models. Current methods perturb the dataset to mark models trained on it. The main challenge lies in crafting this dataset perturbation to balance two competing objectives. On one hand, the perturbation should not significantly degrade performance, preserving the value of training on the data for authorized parties. On the other hand, the modification should sufficiently alter model behavior to enable high-confidence detection.
@@
These objectives are complemented by the following technical requirements. First, the detection should effective with strong guarantees against false positives (models wrongfully detected), making it challenging for dishonest model owners to claim false detection. Second, the detection should rely on black-box model access, ideally top-$k$ predictions for small $k$, to be applicable to models available only through restricted APIs. Finally, the perturbation should be stealth to prevent easy removal by dishonest users, and robust across different model architectures and training recipes.
Previous works either provide strong theoretical guarantees or allow for black-box assessments, but not both. For instance, Radioactive data (Sablayrolles et al., 2020) emphasized statistical significance testing as the theoretical basis to control false positives. However, it marked the weights of the last layer of image classifiers, which is not directly amenable to black-box testing. Conversely, backdoor watermarking methods (Li et al., 2020; 2022; Wenger et al., 2022; Tang et al., 2023) perturb the dataset to predictably alter model predictions when a trigger pattern $t$ is added to a legitimate image $x$. However, they fail to provide guarantees on a benign model’s prediction on watermarked samples $x+t$ watermarking methods assume the . (Guo et al., 2023)
@@
In this paper, we introduce a novel dataset ownership verification approach that enables black-box detection of dishonest models with rigorous theoretical guarantees. We call this method data taggants by analogy to taggants, physical or chemical markers added to materials to trace their usage or manipulation. Primarily designed for image classification datasets, similarly to most prior works (e.g., Sablayrolles et al. (2020); Li et al. (2020); Tang et al. (2023)), the main contributions of data taggants are the following:
  • We transition from triggers used in backdoor watermarking, to keys. A key is a (pattern, label) pair $(x^{(key)},y^{(key)})$. Our method adds a “signature” to the dataset (the taggants) to enforce $f(x^{(key)})\approx y^{(key)}$ for every key when model $f$ has been trained on the signed data. [^1] In the literature on backdoor watermarking such as Souri et al. (2021), triggers are added to legitimate images. A trigger is a pattern $x^{(trigger)}$ that comes together with a target label $y^{(trigger)}$ and the dataset is altered with the goal of enforcing $f(x^{(legit)}+x^{(trigger)})\approx y^{(trigger)}$ for every legitimate image $x^{(legit)}$.
  • While previous work use a single trigger pattern, we use multiple keys, with key labels chosen uniformly at random among possible classes. This randomness enables the use of Binomial significance tests, by checking the proportion of keys for which the top-$k$ predictions contain the key label in response to key patterns, which only requires black-box access to a suspicious model.
Technically, we induce the intended model behavior on the keys using gradient alignment techniques from clean-label data poisoning (Geiping et al., 2020) augmented with a perceptual loss function. This ensures the stealthiness of the taggants and their effectiveness. As we show in our experiments, this approach is more effective than adding the key patterns in transparency on legitimate images.
We perform experiments on ImageNet1k with vision transformers and ResNet architectures of different sizes, together with state-of-the-art training recipes (Wightman et al., 2021) including the 3-Augment data augmentation techniques (Touvron et al., 2022). Our main results are:
  • Under similar conditions for training the dishonest model and for crafting the data taggants, our approach detects dishonest models with high confidence, given access to only top-$10$ predictions.
  • We present a robustness “stress test” through an extensive evaluation of taggants by varying the type of data augmentation used, as well as model architectures and sizes, both for crafting the taggants and for the dishonest model. Our results show that our approach is robust to these variations, although overall achieving lower detection confidence than when the conditions are the same.
  • We explore the stealthiness of taggants using PSNR, out-of-distribution (OOD) detection tests. Additionally, by visual inspection, we observe that the perceptual loss makes taggants less perceptible. Our results also highlight the weakness of OOD approaches based on clustering to detect taggants.
@@ 2 Related work
Steganography is the practice of concealing a message within content, while watermarking is the process of marking data with a message related to that data (often as a proof of ownership) (Cox et al., 2007). They intersect in the contexts where watermarks strive to be imperceptible to avoid content degradation Kahng et al. (1998). Recent works have shown that deep learning methods can improve watermarking technology (Vukotić et al., 2018; Zhu et al., 2018; Fernandez et al., 2022). Furthermore, methods have been developed to watermark both datasets (Guo et al., 2023) and AI models (Adi et al., 2018; Fernandez et al., 2023). While data watermarks are proofs of ownership like our data taggants, watermarks are not designed to persist through processes that use the data. In contrast, data taggants leave a mark on models trained on the dataset, in addition, that mark can be retrieved with only black-box access to the model.
@@
Our work, similarly to most works on dataset ownership verification, focuses on image classification datasets as their main use case. Therefore, we focus the discussion on this case. An early approach to dataset ownership verification involved the modification of certain images in the dataset to align the last activation layer of a classifier with a random direction (Sablayrolles et al., 2020). This method demonstrated success in terms of stealthiness and provided robust theoretical guarantees in a white-box scenario, where access to the weights of the suspicious model is available. However, in the more realistic black-box scenario, the authors only proposed an indirect approach involving learning a copy of the suspicious model by distillation, which requires a high number of queries. More recently, attention has shifted towards backdoor watermarking for dataset ownership verification (Li et al., 2020; 2022; Wenger et al., 2022; Tang et al., 2023; Guo et al., 2023). These methods are closer to our black-box approach, although they rely on confidence scores (Li et al., 2020; 2022) rather than top-k predictions. The data is manipulated such that models trained on it alter their confidence scores when a trigger pattern is added to an input image. From a theoretical perspective, these approaches currently lack robust statistical tests. Some authors have proposed the computation of p-values (Li et al., 2022; Tang et al., 2023), but there is no proof that an honest model satisfies their null hypothesis – that a model not trained on the data should not change its confidence scores or predictions less than a predetermined threshold. The validity of the test is thus questionable, especially since they do not provide theoretical or empirical support for the choice of the threshold on confidence scores. In contrast, our approach enables the application of standard statistical tests.
@@
Our research, akin to prior backdoor watermarking studies, repurposes data poisoning techniques for dataset ownership verification. Data poisoning originally investigates how subtle changes to training data can compromise a model by malicious actors. Two main types of data poisoning attacks exist: backdoor attacks, which modify the model’s behavior when a specific trigger is applied to a class of data (Li et al., 2019; Souri et al., 2021), and targeted attacks, which induce errors on a specific set of inputs (Shafahi et al., 2018; Geiping et al., 2020). These, in turn, may add incorrectly labeled data Li et al. (2020; 2022) to the dataset, which impedes the stealthiness of the approach, or use clean-label modifications (Geiping et al., 2020). Previous studies on backdoor watermarking were mostly based on backdoor attacks, attempting to predictably degrade performance when a trigger is introduced to the data. In contrast, we aim to design models that predict specific key labels in response to key patterns, which is more similar to targeted poisoning attacks – the difference with a poisoning attack being that we alter the behavior on randomly generated patterns rather than legitimate data. Our algorithmic approach leverages gradient matching techniques developed in the clean-label data poisoning literature, for both targeted and backdoor attacks (Geiping et al., 2020; Souri et al., 2021).
@@
The goal of membership inference attacks is to reveal confidential information by inferring which data points were in the training set, usually recognized by low-loss inputs (Shokri et al., 2017; Watson et al., 2021). These methods do not offer any theoretical membership certificate, since a model might have low loss on a sample regardless of whether this sample was actually part of the training set. They are therefore not applicable to dataset ownership verification.
@@ 3 Data Taggants
In the application scenario we consider, Alice wants to publish online a dataset $\mathcal{D}_{A}$. Alice suspects that once $\mathcal{D}_{A}$ is published, Bob will train a model $\mathcal{M}^{B}$ on $\mathcal{D}_{A}$. Given black-box access (top-$k$ predictions) to Bob’s (future) model $\mathcal{M}^{B}$, Alice wants to mark her dataset in order to determine if $\mathcal{M}^{B}$ has been trained on $\mathcal{D}_{A}$.
We propose a solution to Alice’s problem called data taggants, which alters the dataset to mark models trained on it. Data taggants uses data poisonning to induce a certain behavior on Bob’s model, and statistical tests to detect if Bob’s model displays said behavior. To ensure stealthiness, we take inspiration from a clean-label data poisoning, which leaves labels untouched (Geiping et al., 2020).
3.1 Overview
Let us denote Alice’s original dataset by $\mathcal{D}_{A}=(x_{i},y_{i})_{i=1}^{N}\in(\mathcal{X}\times\mathcal{Y})^{N}$ , where $N$ is the number of samples, $\mathcal{X}$ is the set of images and $\mathcal{Y}$ is the set of possible labels. The process of adding data taggants is the following (Figure 1):
@@ 3.1 Overview
    Alice generates a set of $K$ secret keys: $\mathcal{D}_{(key)}=(x^{(key)}_{i},y^{(key)}_{i})_{i=1}^{K}\in(\mathcal{X}\times\mathcal{Y})^{K}$;
@@
    Alice signs her dataset by perturbating the images in a small subset $\mathcal{D}_{S}\subseteq\mathcal{D}_{A}$ of size $S$ called the signing set. The perturbation of images in $\mathcal{D}_{S}$ is called the signature, while data taggants refer to the modified signing set.
@@
    Alice shares a dataset $\tilde{\mathcal{D}}_{A}$, equal to $\mathcal{D}_{A}$ with the modified $\mathcal{D}_{S}$.
The goal of the taggants is to have models trained on $\tilde{\mathcal{D}}_{A}$ predict $y^{(key)}_{i}$ in response to $x^{(key)}_{i}$, while being stealth so that Bob cannot easily remove it. We first formalize Alice’s goal in an ideal scenario where she knows in advance the details of Bob’s training pipeline and discuss our more practical solution afterwards. Let us denote by $\mathcal{L}^{B}_{\theta}$ the loss of Bob’s model with parameter $\theta$. We use $t$ to denote a data augmentation sampled according to Bob’s recipe and $\mathbb{E}_{t}[.]$ the expectation over this random sampling. Alice defines a constrained set of image perturbations $\mathcal{C}=\{\delta\in\mathbb{R}^{d}/\|\delta\|_{\infty}\leq\varepsilon\}$, where $\varepsilon>0$ is kept small to ensure stealthiness. The space of possible signatures is then $\mathcal{C}_{\mathcal{S}}=\{\Delta=(\delta_{j})_{j=1}^{N}\in\mathcal{C}^{N}|\forall j\notin\mathcal{D}_{S},\delta_{j}=0\}$. Alice aims to find the signature $\Delta$ which minimizes the loss on secret keys of Bob’s model after training on $\tilde{\mathcal{D}}_{A}$. This corresponds to the following bilevel optimization problem:
@@
Since solving the bilevel optimization problem above is too costly in terms of compute, we use a variant of gradient matching (Geiping et al., 2020), where the goal is to find $\Delta$ such that, when $\theta\approx\theta^{*}(\Delta)$, $\frac{1}{K}\sum_{i=1}^{K}\nabla_{\theta}\mathcal{L}^{B}_{\theta}(x^{(key)}_{i},y^{(key)}_{i})\approx\frac{1}{S}\sum_{j\in\mathcal{D}_{S}}\mathbb{E}_{t}\big{[}\nabla_{\theta}\mathcal{L}^{B}_{\theta}(t(x_{j}+\delta_{j}),y_{j})\big{]}.$
In reality, Alice does not know in advance Bob’s training algorithm. There may also be multiple Bobs using different training recipes. Our approach involves a model and a set of data augmentation that Alice uses as a surrogate to Bob’s training pipeline. We later study, in our experiments, the robustness of our approach to variations in Alice and Bob’s models and data augmentations.
3.2 Signing the data
The starting point of our approach is that Alice trains a model on her original dataset. We now give the details of the various steps to craft the taggants. We denote by $\theta^{*}$ the parameters of Alice’s model, and $\mathcal{L}^{A}_{\theta^{*}}$ its loss function.

### v2

@@ Abstract
Dataset ownership verification, the process of determining if a dataset is used in a model’s training data, is necessary for detecting unauthorized data usage and data contamination. Existing approaches, such as backdoor watermarking, rely on inducing a detectable behavior into the trained model on a part of the data distribution. However, these approaches have limitations, as they can be harmful to the model’s performances or require unpractical access to the model’s internals. Most importantly, previous approaches lack guarantee against false positives.
This paper introduces data taggants, a novel non-backdoor dataset ownership verification technique. Our method uses pairs of out-of-distribution samples and random labels as secret keys, and leverages clean-label targeted data poisoning to subtly alter a dataset, so that models trained on it respond to the key samples with the corresponding key labels. The keys are built as to allow for statistical certificates with black-box access only to the model.
We validate our approach through comprehensive and realistic experiments on ImageNet1k using ViT and ResNet models with state-of-the-art training recipes. Our findings demonstrate that data taggants can reliably detect models trained on the protected dataset with high confidence, without compromising validation accuracy, and show their superiority over backdoor watermarking. We demonstrate the stealthiness and robustness of our method against various defense mechanisms.
@@ 1 Introduction
Dataset ownership verification (DOV) approaches aim to equip dataset owners with the ability to track the usage of their data in specific trained models. Current methods perturb the dataset to mark models trained on it and the main challenge lies in crafting this dataset perturbation to balance two competing objectives. On the one hand, the perturbation should not significantly degrade performance, preserving the value of training on the data for authorized parties. On the other hand, the modification should sufficiently alter models’ behaviors to enable high-confidence detection.
@@
These objectives are complemented by the following technical requirements. First, the detection should be effective with strong guarantees against false positives (i.e. models wrongfully detected), making it challenging for dishonest model owners to claim false detection. Second, the perturbation should be stealthy to prevent easy removal by dishonest users. Third, it should be robust across different model architectures and training recipes, as to allow for usage in the wild, and adapt to the diversity of models and learning algorithms. Finally, for our method to be practical, the detection should be possible with black-box access to the model. Ideally it should work with top-$k$ predictions for small $k$, to be applicable to models available only through restricted APIs.
Previous works either did not provide strong theoretical guarantees (Maini, 2021; Li et al., 2022; Wenger et al., 2022), did not yield stealthy enough watermarks (Li et al., 2020b), did not demonstrate robustness of the watermark (Li et al., 2020a; Maini, 2021) or only partially allowed for black-box detection (Sablayrolles et al., 2020). Backdoor watermarking, predominantly studied in DOV literature, perturbs the dataset to predictably alter model predictions $f$ when a trigger pattern $x^{(trigger)}$ is added to a legitimate image $x$. This trigger should steer the prediction on the triggered sample away from the ground truth class $y$ and towards a target class $y^{(trigger)}$. The detection of suspicious models is done by running a statistical test to measure a difference between the benign prediction $f(x)_{y}$ and the triggered prediction $f(x+x^{(trigger)})_{y}$. This approach not only contradicts adversarial robustness (a desirable property for deep learning models), but is also harmful to the model as it introduces errors (Guo et al., 2023). More importantly, without proper characterization of a benign model’s predictions, previous detection scheme lack theoretical grounding.
@@
In this paper, we introduce a novel dataset ownership verification approach that enables black-box detection of dishonest models with rigorous theoretical guarantees. We call this method data taggants by analogy to taggants, physical or chemical markers added to materials to trace their usage or manipulation. We introduce a new detection approach: keys, an out-of-distribution (pattern, label) pair $(x^{(key)},y^{(key)})$. When a model $f$ is trained on the protected dataset, it should predict $f(x^{(key)})=y^{(key)}$ for every key. By targeting out-of-distribution key patterns $x^{(key)}$, for which a natural behavior is undefined, we limit the possibilities of inducing errors in the model contrary to backdoor watermarking. This behavior is induced through gradient matching, a clean-label targeted data poisoning technique introduced in Geiping et al. (2020). We perform experiments on ImageNet1k with vision transformers and ResNet architectures of different sizes, together with state-of-the-art training recipes (Wightman et al., 2021) including the 3-Augment data augmentation techniques (Touvron et al., 2022). Primarily designed for image classification datasets, similarly to prior works (Sablayrolles et al., 2020; Li et al., 2022), our main contributions are the following:
  • We introduce a new detection approach: keys, out-of-distribution (pattern, label) pairs $(x^{(key)},y^{(key)})$ on which we measure the top-$k$ accuracy, with $y^{(key)}$ randomly chosen.
  • This randomness enables independence in the use of a theoretically more grounded Binomial significance tests, compared to previous work’s use of pair-wise t-test.
  • We demonstrate the effectiveness and practicality of data taggants through extensive experiments on ImageNet1k with state of the art training procedures. We show the robustness of data taggants when transferred on various model architectures and training recipes. We also bring evidence of the stealthiness of data taggants through PSNR, out-of-distribution (OOD) detection tests, and data poisoning defense approaches. All of this is achieved by modifying only 0.1% of the data and without degradation of performance.
  • We introduce the use of a perceptual loss in the crafting of data poisons to enhance stealthiness and show it allows for visually imperceptible data taggants.
@@ 2 Related work
Steganography is the practice of concealing a message within content, while watermarking is the process of marking data with a message related to that data (often as a proof of ownership) (Cox et al., 2007). They intersect in the contexts where watermarks strive to be imperceptible to avoid content degradation (Kahng et al., 1998). Recent works have shown that deep learning can improve watermarking technology (Vukotić et al., 2018; Zhu et al., 2018; Fernandez et al., 2022) and conversely, AI models can be watermarked (Adi et al., 2018). Watermarks are designed to be detected in data and their radioactivity on models (Sablayrolles et al., 2020; Sander et al., 2024) is only a byproduct. In contrast, data taggants are designed not to be detected, but to leave a mark on models that can later be detected with high confidence.
@@
Our work, similarly to most works on dataset ownership verification (DOV), focuses on image classification datasets as their main use case. Therefore, we focus the discussion on this case. An early approach to DOV involved the modification of certain images in the dataset to align the last activation layer of a classifier with a random direction (Sablayrolles et al., 2020). This method demonstrated success in terms of stealthiness and provided robust theoretical guarantees in a white-box scenario, where access to the weights of the suspicious model is available. However, in the more realistic black-box scenario, the authors only proposed an indirect approach involving distillating the suspicious model, which requires a high number of queries  [^1] Black-box without distillation version of radioactive data amounts to a membership inference attack which hence lack the strong theoretical guarantees of the approach.. More recently, attention has shifted towards backdoor watermarking for DOV (Li et al., 2020a; 2022; Wenger et al., 2022; Tang et al., 2023; Guo et al., 2023). These methods are closer to our black-box approach, although they require more information from the model and rely on confidence scores (Li et al., 2020a; 2022) rather than top-k predictions. The data is manipulated such that models trained on it alter their confidence scores when a trigger pattern is added to an input image. From a theoretical perspective, these approaches currently lack theoretical grounding. They all rely on the assumption that an honest model satisfies their null hypothesis – i.e. that a model not trained on the data should not change its confidence scores or predictions more than a predetermined threshold. The validity of the test is thus questionable, especially since they do not provide theoretical or empirical support for the choice of the threshold on confidence scores. In contrast, our approach precisely characterize the behavior of a benign model which enables the application of standard and sound statistical tests. Table 5 in Section A.1 in Appendix compares our work with prior approaches.
@@
Our research, akin to prior backdoor watermarking studies, repurposes data poisoning techniques for DOV. Data poisoning originally investigates how subtle changes to training data can compromise a model by malicious actors. Two main types of data poisoning attacks exist: backdoor attacks, which modify the model’s behavior when a specific trigger is applied to a class of data (Li et al., 2019; Souri et al., 2021), and targeted attacks, which induce errors on a specific set of inputs (Shafahi et al., 2018; Geiping et al., 2020). These, in turn, may add incorrectly labeled data Li et al. (2020a; 2022) to the dataset, which impedes the stealthiness of the approach. Previous studies on backdoor watermarking were mostly based on backdoor attacks, attempting to predictably degrade performance when a trigger is introduced to the data. In contrast, we build on top of a clean-label targeted data poisoning approach (Geiping et al., 2020). We aim to design models that predict specific key labels in response to key inputs. The difference with a poisoning attack being that we alter the behavior on randomly generated patterns rather than legitimate data as to not induce malicious errors. We refer to that approach as data signing. Our algorithmic approach leverages gradient matching techniques developed in the clean-label data poisoning literature, for both targeted and backdoor attacks (Geiping et al., 2020; Souri et al., 2021). This technique allows to reproduce, via data poisoning, the effects of a particular gradient direction, which has been shown to be possible even for arbitrary gradient attacks Bouaziz et al. (2024).
@@
The goal of membership inference attacks is to reveal confidential information by inferring which data points were in the training set, usually recognized by low-loss inputs (Shokri et al., 2017; Watson et al., 2021). These methods do not offer any theoretical membership certificate, since a model might have low loss on a sample for different reasons than this sample simply being in the training set. MIAs are not applicable to DOV (Zhang et al., 2024).
@@ 3 Data Taggants
In the application scenario we consider, Alice wants to publish online a dataset $\mathcal{D}_{A}$. Alice suspects Bob will try to train a model $\mathcal{M}^{B}$ on $\mathcal{D}_{A}$. Given black-box access (top-$k$ predictions) to Bob’s model $\mathcal{M}^{B}$, Alice wants to mark her dataset in order to determine if $\mathcal{M}^{B}$ was trained on $\mathcal{D}_{A}$.
We propose a solution to Alice’s problem called data taggants, which alters the dataset to mark models trained on it. Data taggants uses data poisoning to induce a certain behavior on Bob’s model, and statistical tests to detect if Bob’s model displays said behavior. To ensure stealthiness, we take inspiration from clean-label data poisoning, which leaves labels untouched (Geiping et al., 2020). Since the goal of our approach is to harmlessly influence the model, we designate our approach as data signing instead of data poisoning.
3.1 Overview and technical background
Let us denote Alice’s original dataset by $\mathcal{D}_{A}=\{(x_{i},y_{i})_{i=1}^{N}\}\in(\mathcal{X}\times\mathcal{Y})^{N}$ , where $N$ is the number of samples, $\mathcal{X}$ is the input space and $\mathcal{Y}$ is the set of possible labels. The process of adding data taggants is the following (Figure 1):
@@ 3.1 Overview and technical background
    Alice generates a set of $K$ secret keys: $\mathcal{D}_{(key)}=\{(x^{(key)}_{i},y^{(key)}_{i})_{i=1}^{K}\}\in(\mathcal{X}\times\mathcal{Y})^{K}$;
@@
    Alice signs (i.e. harmlessly poisons) her dataset by perturbating the images in a small subset $\mathcal{D}_{S}=\{(x^{(sign)}_{i},y^{(sign)}_{i})_{i=1}^{K}\}\subseteq\mathcal{D}_{A}$ of size $S$ called the signing set. The perturbations $\Delta=\{\delta_{i},i\in[S]\}$ added on top of images in $\mathcal{D}_{S}$ are called signatures, while data taggants refer to the modified signing set $x^{(taggant)}_{i}=x^{(sign)}_{i}+\delta_{i},\forall i\in[S]$.
@@
    Alice shares $\tilde{\mathcal{D}}_{A}$, a modified version of $\mathcal{D}_{A}$ with the crafted data taggants replacing the elements in the signing set $\mathcal{D}_{S}$. The keys are kept secret and never shared with Bob.
The goal of the data taggants is to have models trained on $\tilde{\mathcal{D}}_{A}$ predict $y^{(key)}_{i}$ in response to $x^{(key)}_{i}$, while being stealthy so that Bob cannot easily remove it, and robust to different settings (model architecture, training algorithm) Bob could use. Let us denote by $\mathcal{L}^{B}_{\theta}$ the loss of Bob’s model with parameter $\theta$. We use $t$ to denote a data augmentation sampled according to Bob’s recipe and $\mathbb{E}_{t}[\cdot]$ the expectation over this random sampling. Alice defines a constrained set of image perturbations $\mathcal{C}=\{\delta\in\mathbb{R}^{d}/\|\delta\|_{\infty}\leq\varepsilon\}$, where $\varepsilon>0$ is kept small to ensure stealthiness. The space of possible signatures is then $\mathcal{C}_{\mathcal{S}}=\{\Delta=(\delta_{j})_{j=1}^{N}\in\mathcal{C}^{N}|\forall j\notin\mathcal{D}_{S},\delta_{j}=0\}$. Alice aims to find the signature $\Delta$ which minimizes the loss of Bob’s model on the keys after training on $\tilde{\mathcal{D}}_{A}$. This corresponds to the following bilevel optimization problem:
@@
Since solving the bilevel optimization problem above is intractable, we use a variant of gradient matching (Geiping et al., 2020), where the goal is to find $\Delta$ such that, when $\theta\approx\theta^{*}(\Delta)$,
$\frac{1}{K}\sum_{i=1}^{K}\nabla_{\theta}\mathcal{L}^{B}_{\theta}(x^{(key)}_{i},y^{(key)}_{i})\approx\frac{1}{S}\sum_{j=1}^{S}\mathbb{E}_{t}\big[\nabla_{\theta}\mathcal{L}^{B}_{\theta}(t(x^{(sign)}_{j}+\delta_{j}),y^{(sign)}_{j})\big$].
So that an optimization step on the data taggants also improves the model’s loss on the keys. In reality, Alice does not know in advance Bob’s training algorithm. There may also be multiple “Bobs” using different model architecture or training recipes. Alice thus cannot anticipate which architecture or training recipe to use to craft data taggants. Our approach involves a model and a set of data augmentation that Alice uses as a surrogate to Bob’s training pipeline. We later study, in our experiments, the robustness of our approach when Alice and Bob use different model architectures or sets of data augmentations.
3.2 Signing the data in practice
The starting point of our approach is that Alice trains a model on her original dataset. We denote by $\theta^{*}$ the parameters of Alice’s model, and $\mathcal{L}^{A}_{\theta^{*}}$ its loss function. We now give the details of the various steps to craft the taggants.